DeLong Corporation was organized on January 1, 2019. It is authorized to issue 10,000 shares...


P11-1A DeLong Corporation was organized on January 1, 2019. It is authorized to issue 10,000 shares of 8%, $100 par value preferred stock, and 500,000 shares of no-par common stock with a stated value of $2 per share. The following stock transactions were completed during the first year. 


Jan. 10 Issued 80,000 shares of common stock for cash at $4 per share. 

Mar. 1 Issued 5,000 shares of preferred stock for cash at $105 per share. 

Apr. 1 Issued 24,000 shares of common stock for land. The asking price of the land was $90,000. The fair value of the land was $85,000. 

May. 1 Issued 80,000 shares of common stock for cash at $4.50 per share. 

Aug 1 Issued 10,000 shares of common stock to attorneys in payment of their bill of $30,000 for services performed in helping the company organize. 

Sept 1 Issued 10,000 shares of common stock for cash at $5 per share. 

Nov. 1 Issued 1,000 shares of preferred stock for cash at $109 per share.. 


Instructions 

(a) Journalize the transactions. 

(b) Post to the stockholders' equity accounts. (Use T-accounts.) 

(c) Prepare the paid-in capital section of stockholders' equity at December 31, 2019.

 (c) Total paid-in capital $1,479,000 

Journalize and post treasury stock transactions, and prepare stockholders' equity section. 
Answer #1


DeLong Corporation


Journal entries


DateAccountDebitCreditCalculation
10-JanCash$ 320,000
=80,000*4

Common Stock
$   160,000=80,000*2

Paid in Capital in Excess of Stated Value-Common Stock
$   160,000





1-MarCash$ 525,000
=5,000*105

Preferred Stock
$   500,000=5,000*100

Paid in Capital in Excess of Par-Preferred Stock
$     25,000





1-AprLand$    85,000


Common Stock
$     48,000=24,000*2

Paid in Capital in Excess of Stated Value-Common Stock
$     37,000





1-MayCash$ 360,000
=80,000*4.5

Common Stock
$   160,000=80,000*2

Paid in Capital in Excess of Stated Value-Common Stock
$   200,000





1-AugOrganization expenses$    30,000


Common Stock
$     20,000=10,000*2

Paid in Capital in Excess of Stated Value-Common Stock
$     10,000





1-SepCash$    50,000
=10,000*5

Common Stock
$     20,000=10,000*2

Paid in Capital in Excess of Stated Value-Common Stock
$     30,000





1-NovCash$ 109,000
=1,000*109

Preferred Stock
$   100,000=1,000*100

Paid in Capital in Excess of Par-Preferred Stock
$        9,000

DeLong Corporation

Balance Sheet Partial

As of December 31, 2019




Stockholders Equity

Capital Stock

Preferred Stock         600,000
Common Stock         408,000
Total Capital Stock
     1,008,000
Additional Paid in capital

Paid in Capital in Excess of Par-Preferred Stock           34,000
Paid in Capital in Excess of Stated Value-Common Stock         437,000
Total Additional Paid in Capital
         471,000
Total Stockholders Equity
     1,479,000
Answer #2
DeLong Corporation
Journal entries
Date Account Debit Credit Calculation
10-Jan Cash $ 320,000 =80,000*4
Common Stock $   160,000 =80,000*2
Paid in Capital in Excess of Stated Value-Common Stock $   160,000
1-Mar Cash $ 525,000 =5,000*105
Preferred Stock $   500,000 =5,000*100
Paid in Capital in Excess of Par-Preferred Stock $     25,000
1-Apr Land $    85,000
Common Stock $     48,000 =24,000*2
Paid in Capital in Excess of Stated Value-Common Stock $     37,000
1-May Cash $ 360,000 =80,000*4.5
Common Stock $   160,000 =80,000*2
Paid in Capital in Excess of Stated Value-Common Stock $   200,000
1-Aug Organization expenses $    30,000
Common Stock $     20,000 =10,000*2
Paid in Capital in Excess of Stated Value-Common Stock $     10,000
1-Sep Cash $    50,000 =10,000*5
Common Stock $     20,000 =10,000*2
Paid in Capital in Excess of Stated Value-Common Stock $     30,000
1-Nov Cash $ 109,000 =1,000*109
Preferred Stock $   100,000 =1,000*100
Paid in Capital in Excess of Par-Preferred Stock $        9,000

Date DeLong Corporation T Accounts Date Debit Date Credit Preferred Stock 1-Mar $ 500,000 1-Nov $ 100,000 Debit Date Credit C

DeLong Corporation
Balance Sheet Partial
As of December 31, 2019
Stockholders Equity
Capital Stock
Preferred Stock          600,000
Common Stock          408,000
Total Capital Stock      1,008,000
Additional Paid in capital
Paid in Capital in Excess of Par-Preferred Stock            34,000
Paid in Capital in Excess of Stated Value-Common Stock          437,000
Total Additional Paid in Capital          471,000
Total Stockholders Equity      1,479,000